katherine:

Determine the equation of the line with slope −4 that passes through the point M(−2, 1). y = −4x + 7 y = −4x − 7 y = 4x + 9 y = −4x − 9

katherine:

@Hero @JustSaiyan @umm @Ultrilliam

Hero:

The equation of a line is y = mx + b right?

katherine:

Yes!

Hero:

And point M is (x,y) = (-2,1) and the slope is m = -4 correct?

katherine:

Correct

Hero:

So in other words, x = -2, y = 1 and m = -4. Do you think you might be able to enter those values in to the equation of the line y = mx + b?

katherine:

yea, that's my problem. idk how lmao

Hero:

So for y = mx + b, you don't know how to replace x with -2, replace y with 1 and replace m with -4?

katherine:

i'm sorry. I've been working on this test for 3 hrs and my brain is dying lol

Hero:

Oh, you're taking a test. Well, you should take a look at our rules: http://www.questioncove.com/rules

katherine:

ya.... everyone does it so..